OpenStudy (anonymous):

is integral from 1 to inf of e^x^2 convergent or divergent

zepdrix (zepdrix):

Hmmm...

zepdrix (zepdrix):

I remember you can do something like this...

zepdrix (zepdrix):

\[\Large\rm \int\limits_1^{\infty} e^{x^2}~dx \cdot \int\limits_1^{\infty} e^{y^2}~dx =\int\limits \int\limits e^{x^2+y^2}~dx~dy\]Changing to polar,\[\Large\rm x=r\cos \theta\]\[\Large\rm y=r \sin \theta\]Changing our integral to,\[\Large\rm \int\limits_{\theta}\int\limits_{r} e^{r^2}(r dr~d \theta)\]Making it a lot easier to integrate. Buuuuuut the boundaries, hmmm....

zepdrix (zepdrix):

Oh oh and keep in mind that I multiplied two of the initial integrals together. So if we're calling that integral \(\Large\rm I\), then the first step was \(\Large\rm I^2\). So after doing all the integration, we would need to take the square root of the result.

OpenStudy (anonymous):

hmmm we havn't learned polar coordinates yet... -___- lol is there another way?

ganeshie8 (ganeshie8):

you may use comparison test, look for some converging function f(x) thats greater than e^x^2 in the given integral

OpenStudy (anonymous):

what could my comparison function be?

OpenStudy (anonymous):

simply e^x?

ganeshie8 (ganeshie8):

Notice that \(\large e^{x^2}\) shoots up quickly and so diverges for sure, are you really sure you're not missing a negative sign in the exponent ?

zepdrix (zepdrix):

Oh my bad :( They didn't want the actual integral value, they just wanted convergent divergent :3 hehe woops.

OpenStudy (anonymous):

loll its ok i appreciate all that time you spent making such nice equations :D

OpenStudy (anonymous):

i am not missing anything, although our professor has been known to accidentally give us problems that are too difficult and he might change them when someone askes in class (he writes all his own problems -__-)

ganeshie8 (ganeshie8):

if its really e^x^2, then its easy.. just use e^x as your comparison function

OpenStudy (anonymous):

i dont know if it diverges for sure, i would have that gabriels horn was divergent by looking at the graph but it equals pi -__-

OpenStudy (anonymous):

really? ok and then use LCTT?

ganeshie8 (ganeshie8):

you may want to show the work for evaluating the integral e^x

OpenStudy (anonymous):

isn't e^x^2 bigger than e^x?

ganeshie8 (ganeshie8):

yeah it was a typo sorry to confuse u ;p

OpenStudy (anonymous):

loll ok ;p thank you!

OpenStudy (anonymous):

yay so direct comparison works!

OpenStudy (anonymous):

thank you both :D

ganeshie8 (ganeshie8):

corrected typo : \(\large e^{x^2} \gt e^x\) in the given interval since the smaller function \(\large \int\limits_{1}^{\infty}e^x dx\) diverges, the integral in question also diverges

OpenStudy (anonymous):

k thank you !